# Regularity and stability in a strongly degenerate nonlinear diffusion and haptotaxis model of cancer invasion

### v1

@@ 5 Stability with respect to initial perturbations
$-\int_{0}^{T}\int_{\Omega}\psi\partial_{t}\chi\,{\rm d}{\bm{x}}{\rm d}t=\int_{\Omega}\psi^{0}\chi(0,{\bm{x}})\,{\rm d}{\bm{x}}+\int_{0}^{T}\int_{\Omega}\psi\nabla\chi\cdot{{\bm{V}}}(\psi,\varphi)\,{\rm d}{\bm{x}}{\rm d}t+\int_{0}^{T}\int_{\Omega}\chi\psi\bar{R}(\psi,\varphi)\,{\rm d}{\bm{x}}{\rm d}t$
[…]
@@
It is necessary to introduce a possible weak limit $\bar{R}$ because of the degeneracy of Eq. (1). For example, if $\varphi^{0}\equiv 0$, then initial oscillations in $\psi^{0}_{\varepsilon}$ are not regularized and the weak limit of $\psi_{\varepsilon}R(\psi_{\varepsilon},\varphi^{0})=-R_{0}\psi_{\varepsilon}^{2}$ will nor be $-R_{0}\psi^{2}$.
@@
Notice that, by weak-strong convergence, we know indeed that the limit is $\varphi^{\frac{1}{2}}\psi$. Hence, in the set $\{\varphi(t,{\bm{x}})>0\}$ we also have the strong convergence
$\psi_{\varepsilon}(t,{\bm{x}})\to\psi(t,{\bm{x}})\quad\text{strongly in}\quad L^{p}((0,T)\times B\cap\{\varphi>0\}),\quad 1\leq p<3\,,$
again, by extraction of a subsequence that converges almost everywhere. Finally, if $0<\varphi^{0}({\bm{x}})$, then from Eq. (1)2 (which holds in the limit again by weak-strong convergence, see Step 2) we see that we have $\varphi(t,{\bm{x}})>0$ for all $t\in[0,T]$ and thus conclude that $\psi_{\varepsilon}$ converges almost everywhere and
@@
$-\int_{0}^{T}\int_{\Omega}\psi_{\varepsilon}\partial_{t}\chi\,{\rm d}{\bm{x}}{\rm d}t=\int_{\Omega}\psi_{\varepsilon}^{0}\chi(0,{\bm{x}})\,{\rm d}{\bm{x}}+\int_{0}^{T}\int_{\Omega}\psi_{\varepsilon}\nabla\chi\cdot{{\bm{V}}}(\psi_{\varepsilon},\varphi_{\varepsilon})\,{\rm d}{\bm{x}}{\rm d}t+\int_{0}^{T}\int_{\Omega}\chi\psi_{\varepsilon}R(\psi_{\varepsilon},\varphi_{\varepsilon})\,{\rm d}{\bm{x}}{\rm d}t$ (23)
@@
[…]
@@
$\int_{0}^{T}\int_{\Omega}\psi_{\varepsilon}\nabla\chi\cdot{{\bm{V}}}(\psi_{\varepsilon},\varphi_{\varepsilon})\,{\rm d}{\bm{x}}{\rm d}t=\int_{0}^{T}\int_{\Omega}\alpha\underbrace{\psi_{\varepsilon}\varphi_{\varepsilon}(1-\varphi_{\varepsilon})\;\nabla\chi\cdot\nabla\varphi_{\varepsilon}}_{\mathcal{A}_{\varepsilon}}-D\underbrace{\psi_{\varepsilon}\varphi_{\varepsilon}(1-\varphi_{\varepsilon})\;\nabla\chi\cdot\nabla\psi_{\varepsilon}}_{\mathcal{B}_{\varepsilon}}\,{\rm d}{\bm{x}}{\rm d}t\,.$
From the convergence results detailed in Step 1, the term $\mathcal{A}_{\varepsilon}$ converges strongly to $\mathcal{A}=\psi\,\varphi(1-\varphi)\;\nabla\chi\cdot\nabla\varphi$ because $\psi_{\varepsilon}\varphi_{\varepsilon}(1-\varphi_{\varepsilon})$ converges strongly in $L^{p}((0,T)\times B)$, $31\leq p<$ 3, while $\nabla\varphi_{\varepsilon}$ converges weakly in $L^{2}((0,T)\times B)$. For the other term, we write
@@
The term $(1-\varphi_{\varepsilon})\sqrt{\psi_{\varepsilon}}\varphi_{\varepsilon}^{\frac{1}{4}}$ converges strongly in any $L^{p}((0,T)\times B)$, $61\leq p<$ 6, to $(1-\varphi)\sqrt{\psi}\varphi^{\frac{1}{4}}$ using Step 1, while $\varphi_{\varepsilon}^{\frac{1}{4}}\sqrt{\psi_{\varepsilon}\varphi_{\varepsilon}}\nabla\chi\cdot\nabla\psi_{\varepsilon}$ is bounded in $L^{2}((0,T)\times B)$ thanks to (18), and thus converges weakly. To identify its limit we write it as
@@
The term $\mathcal{C}_{1\varepsilon}$ converges because, as before, $\varphi_{\varepsilon}^{\frac{1}{4}}\sqrt{\psi_{\varepsilon}}$ converges strongly to $\varphi^{\frac{1}{4}}\sqrt{\psi}$ in any $L^{p}((0,T)\times B)$, $61\leq p<$ 6, while $\nabla\chi\cdot\nabla(\sqrt{\varphi_{\varepsilon}}\psi_{\varepsilon})$ converges weakly in $L^{\frac{3}{2}}((0,T)\times B)$ to $\nabla\chi\cdot\nabla(\sqrt{\varphi}\psi)$. The term $\mathcal{C}_{2\varepsilon}$ converges because, as before, $\varphi_{\varepsilon}^{\frac{1}{4}}{\psi_{\varepsilon}}$ converges strongly to $\varphi^{\frac{1}{4}}{\psi}$ in any $L^{p}((0,T)\times B)$, $31\leq p<$ 3, while $\sqrt{\psi_{\varepsilon}}\nabla\chi\cdot\nabla\sqrt{\varphi_{\varepsilon}}$ converges weakly in $L^{2}((0,T)\times B)$ to $\sqrt{\psi}\nabla\chi\cdot\nabla\sqrt{\varphi}$.
@@
Finally, if $0<\varphi^{0}({\bm{x}})$ we have strong convergence of $\psi_{\varepsilon}$ and can pass to the limit also in the second term, concluding that the limit functions satisfy Eq. (1) in the sense of Definition 3, with $\bar{R}\equiv R$ everywhere. Otherwise, we can only conclude this with $\bar{R}=R$ on the set $\{\varphi>0\}$.

### v2

@@ 5 Stability with respect to initial perturbations
$\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi\partial_{t}\chi\,{\rm d}{\bm{x}}{\rm d}t=\int_{\Omega}\psi^{0}\chi(0,{\bm{x}})\,{\rm d}{\bm{x}}+\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi\nabla\chi\cdot{{\bm{V}}}(\psi,\varphi)\,{\rm d}{\bm{x}}{\rm d}t+\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\chi\psi\bar{R}(\psi,\varphi)\,{\rm d}{\bm{x}}{\rm d}t$
$\varphi(t,x)=\varphi^{0}(x)-\gamma\int_{0}^{t}\varphi\psi(s,x)ds,\qquad\forall t\geq 0,\;x\in\Omega,$
@@
It is necessary to introduce a possible weak limit $\bar{R}$ because of the degeneracy of Eq. (1). For example, if $\varphi^{0}\equiv 0$, then initial oscillations in $\psi^{0}_{\varepsilon}$ are not regularized and the weak limit of $\psi_{\varepsilon}R(\psi_{\varepsilon},\varphi^{0})=-R_{0}\psi_{\varepsilon}^{2}$ will not be $R_{0}\psi^{2}$.
@@
Notice that, by weak-strong convergence, we know indeed that the limit is $\varphi^{\frac{1}{2}}\psi$. In the set $\{\varphi(t,{\bm{x}})>0\}$ we know that $\varphi_{\varepsilon}(t,{\bm{x}})\to\varphi(t,{\bm{x}})$ almost everywhere pointwise and thus we also have the strong convergence
$\psi_{\varepsilon}(t,{\bm{x}})\to\psi(t,{\bm{x}})\quad\text{strongly in}\quad L^{p}((0,T)\times B\cap\{\varphi>0\}),\quad 1\leq p<3\,.$
Finally, if $0<\varphi^{0}({\bm{x}})$, then from Eq. (1)2 (which holds in the limit again by weak-strong convergence, see Step 2) we see that we have $\varphi(t,{\bm{x}})>0$ for all $t\in[0,T]$ and thus conclude that $\psi_{\varepsilon}$ converges almost everywhere and
@@
$\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi_{\varepsilon}\partial_{t}\chi\,{\rm d}{\bm{x}}{\rm d}t=\int_{\Omega}\psi_{\varepsilon}^{0}\chi(0,{\bm{x}})\,{\rm d}{\bm{x}}+\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi_{\varepsilon}\nabla\chi\cdot{{\bm{V}}}(\psi_{\varepsilon},\varphi_{\varepsilon})\,{\rm d}{\bm{x}}{\rm d}t+\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\chi\psi_{\varepsilon}R(\psi_{\varepsilon},\varphi_{\varepsilon})\,{\rm d}{\bm{x}}{\rm d}t$ (27)
@@
$\varphi_{\varepsilon}(t,x)=\varphi_{\varepsilon}^{0}(x)-\gamma\int_{0}^{t}\varphi_{\varepsilon}\psi_{\varepsilon}(s,x)ds,\qquad\forall t\geq 0,\;x\in\Omega.$ (28)
@@
$\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi_{\varepsilon}\nabla\chi\cdot{{\bm{V}}}(\psi_{\varepsilon},\varphi_{\varepsilon})\,{\rm d}{\bm{x}}{\rm d}t=\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\alpha\underbrace{\psi_{\varepsilon}\varphi_{\varepsilon}(1-\varphi_{\varepsilon})\;\nabla\chi\cdot\nabla\varphi_{\varepsilon}}_{\mathcal{A}_{\varepsilon}}-D\underbrace{\psi_{\varepsilon}\varphi_{\varepsilon}(1-\varphi_{\varepsilon})\;\nabla\chi\cdot\nabla\psi_{\varepsilon}}_{\mathcal{B}_{\varepsilon}}\,{\rm d}{\bm{x}}{\rm d}t\,.$
From the convergence results detailed in Step 1, the term $\mathcal{A}_{\varepsilon}$ converges strongly to $\mathcal{A}=\psi\,\varphi(1-\varphi)\;\nabla\chi\cdot\nabla\varphi$ because $\psi_{\varepsilon}\varphi_{\varepsilon}(1-\varphi_{\varepsilon})$ converges strongly in $L^{p}((0,T)\times B)$, $1\leq p<3$, while $\nabla\varphi_{\varepsilon}$ converges weakly in $L^{2}((0,T)\times B)$. For the other term, we write
@@
The term $(1-\varphi_{\varepsilon})\sqrt{\psi_{\varepsilon}}\varphi_{\varepsilon}^{\frac{1}{4}}$ converges strongly in any $L^{p}((0,T)\times B)$, $1\leq p<6$, to $(1-\varphi)\sqrt{\psi}\varphi^{\frac{1}{4}}$ using Step 1, while $\varphi_{\varepsilon}^{\frac{1}{4}}\sqrt{\psi_{\varepsilon}\varphi_{\varepsilon}}\nabla\chi\cdot\nabla\psi_{\varepsilon}$ is bounded in $L^{2}((0,T)\times B)$ thanks to (19), and thus converges weakly. To identify its limit we write it as
@@
The term $\mathcal{C}_{1\varepsilon}$ converges because, as before, $\varphi_{\varepsilon}^{\frac{1}{4}}\sqrt{\psi_{\varepsilon}}$ converges strongly to $\varphi^{\frac{1}{4}}\sqrt{\psi}$ in any $L^{p}((0,T)\times B)$, $1\leq p<6$, while $\nabla\chi\cdot\nabla(\sqrt{\varphi_{\varepsilon}}\psi_{\varepsilon})$ converges weakly in $L^{\frac{3}{2}}((0,T)\times B)$ to $\nabla\chi\cdot\nabla(\sqrt{\varphi}\psi)$. The term $\mathcal{C}_{2\varepsilon}$ converges because, as before, $\varphi_{\varepsilon}^{\frac{1}{4}}{\psi_{\varepsilon}}$ converges strongly to $\varphi^{\frac{1}{4}}{\psi}$ in any $L^{p}((0,T)\times B)$, $1\leq p<3$, while $\sqrt{\psi_{\varepsilon}}\nabla\chi\cdot\nabla\sqrt{\varphi_{\varepsilon}}$ converges weakly in $L^{2}((0,T)\times B)$ to $\sqrt{\psi}\nabla\chi\cdot\nabla\sqrt{\varphi}$.
@@
Finally, on the set $\{\varphi>0\}$ we have proved pointwise convergence both for $\varphi_{\varepsilon}$ and $\psi_{\varepsilon}$, therefore $R_{\varepsilon}$ converges a.e. to $R=R_{0}(1-\varphi-\psi)$. So the pointwise limit of $\psi_{\varepsilon}R_{\varepsilon}$ is $\psi R$ and it is also the limit in all $L^{p}$, $p<3$, spaces. Otherwise, we can only conclude that $R_{\varepsilon}$ converges weakly to some function $\bar{R}$, with $\bar{R}=R$ on the set $\{\varphi>0\}$, but we do not know what $\psi_{\varepsilon}^{2}$ converges to on the set $\{\varphi=0\}$. We thus conclude that the limit functions $(\psi,\varphi)$ satisfy Eq. (1) in the sense of Definition 3
6 Existence of solutions
Following the strategy adopted in [37, 35], we infer existence of weak solutions of the degenerate problem from an approximating problem with a non-degenerate velocity field, exploiting the regularity established in Sections 3.1-4.
For $\delta>0$, let $\psi_{\delta}(t,{\bm{x}})\geq 0$ and $\varphi_{\delta}(t,{\bm{x}})\geq 0$ satisfy in the weak sense the system
$\begin{cases}\partial_{t}\psi_{\delta}+\nabla\cdot\left[\psi_{\delta}\big(\delta+\varphi_{\delta}(1-\varphi_{\delta})\big)\big(\alpha\nabla\varphi_{\delta}-D\nabla\psi_{\delta}\big)\right]=\psi_{\delta}\,R_{0}(1-\varphi_{\delta}-\psi_{\delta}),\\ \partial_{t}\varphi_{\delta}=-\gamma\varphi_{\delta}\psi_{\delta},\end{cases}$ (29)
for $t\in[0,T]$ and ${\bm{x}}\in\Omega\subseteq\mathbb{R}^{d}$ ($d=1,2,3$), complemented with initial conditions $\psi_{\delta}(0,{\bm{x}})=\psi^{0}_{\delta}$ and $\varphi_{\delta}(0,{\bm{x}})=\varphi^{0}_{\delta}$ satisfying the properties (6) and (7). This system is obtained from the degenerate one by relaxing the saturation term (3) mapping $S\to S_{\delta}$ with
$S_{\delta}(\varphi):=\delta+S(\varphi)>0$
for all $\varphi\in[0,1]$ – cf. Figure 1 – or, analogously, $V\to V_{\delta}$ with
$V_{\delta}(\psi,\varphi):=S_{\delta}(\varphi)\,(\alpha\nabla\varphi-D\nabla\psi)\equiv\delta\,(\alpha\nabla\varphi-D\nabla\psi)+V(\psi,\varphi)$
where it is clear that the degeneracy at $\varphi=0$ disappears.
6.1 Regularity results and compactness
Let us extend the results of Sections 3.1-4 to the non-degenerate problem.
Proposition 5 (Bounds and compactness in the non-degenerate problem)
Let $\Omega\equiv\mathbb{R}^{d}$, and let $\psi_{\delta}\geq 0$ and $\varphi_{\delta}\geq 0$ be solutions of (29) complemented with initial conditions $\psi_{\delta}(0,{\bm{x}})=\psi^{0}_{\delta}$ and $\varphi_{\delta}(0,{\bm{x}})=\varphi^{0}_{\delta}$ satisfying assumptions (6) and (7). Then we have that $\psi_{\delta}\geq 0$ and $\varphi_{\delta}\geq 0$ satisfy uniformly in $\delta$ all bounds listed in Propositions 1 and 2, as well as the following bound for all $\delta>0$ and $T>0$:
$\delta\,\psi_{\delta}\,|D\nabla\psi_{\delta}-\alpha\nabla\varphi_{\delta}|^{2}\in L^{1}((0,T)\times\Omega$ (30)
Moreover, we have local compactness of $\varphi_{\delta}$ and $\varphi^{\frac{1}{2}}_{\delta}\psi_{\delta}$ in space and time.
Proof.
Step 1. The bounds in Proposition 1 and the new estimate. The proof of bounds (8), (9) and (11)-(13) follows the same steps detailed in the proof of Proposition 1. Meanwhile the energy has an extra term, yielding
$\begin{split}\frac{d}{dt}\int_{\Omega}\psi\left(\frac{D}{2}\psi_{\delta}-\alpha\varphi_{\delta}\right)\,{\rm d}{\bm{x}}+\delta\int_{\Omega}\psi_{\delta}|D\nabla\psi_{\delta}-\alpha\nabla\varphi_{\delta}|^{2}{\rm d}{\bm{x}}\leq\;&C_{2}\int_{\Omega}\psi_{\delta}\left(\frac{D}{2}\psi_{\delta}-\alpha\varphi_{\delta}\right)\,{\rm d}{\bm{x}}\\ &+\alpha(C_{2}+R_{0})\int_{\Omega}\psi_{\delta}\,{\rm d}{\bm{x}},\end{split}$
from which we can again obtain (16). Moreover, integrating in time, we conclude
$\delta\int_{0}^{T}\int_{\Omega}\psi_{\delta}\,|D\nabla\psi_{\delta}-\alpha\nabla\varphi_{\delta}|^{2}\,{\rm d}{\bm{x}}\,{\rm d}t\leq C_{12}(T)\,.$
thanks to (6), (15) and (16), i.e. we have the new bound (30). This new estimate allows us to conclude the proof of bound (10). In fact, the second moment also has an additional term
$\frac{d}{dt}\int_{\Omega}\frac{|{\bm{x}}|^{2}}{2}\psi_{\delta}(t,{\bm{x}})\,{\rm d}{\bm{x}}=$ $\int_{\Omega}|{\bm{x}}|\,\psi_{\delta}\varphi_{\delta}(1-\varphi_{\delta})|\alpha\nabla\varphi_{\delta}-D\nabla\psi_{\delta}|\,{\rm d}{\bm{x}}+\int_{\Omega}\frac{|{\bm{x}}|^{2}}{2}\psi_{\delta}R_{0}(1-\psi_{\delta}-\varphi_{\delta})\,{\rm d}{\bm{x}}$
$+\delta\int_{\Omega}|{\bm{x}}|\,\psi_{\delta}\,|D\nabla\psi_{\delta}-\alpha\nabla\varphi_{\delta}|\,{\rm d}{\bm{x}}$
to which we can apply the Cauchy-Schwarz and the AM-GM inequalities, as we did to the first term (see Step 3 in proof of Proposition 1), to obtain
$\frac{d}{dt}\int_{\Omega}\frac{|{\bm{x}}|^{2}}{2}\psi_{\delta}(t,{\bm{x}})\,{\rm d}{\bm{x}}\leq$ $\int_{\Omega}\frac{|{\bm{x}}|^{2}}{2}\psi_{\delta}\varphi_{\delta}(1-\varphi_{\delta})\,{\rm d}{\bm{x}}+\frac{1}{2}\int_{\Omega}\psi_{\delta}\varphi_{\delta}(1-\varphi_{\delta})|\alpha\nabla\varphi_{\delta}-D\nabla\psi_{\delta}|^{2}\,{\rm d}{\bm{x}}$
$+\delta\int_{\Omega}\frac{|{\bm{x}}|^{2}}{2}\psi_{\delta}\,{\rm d}{\bm{x}}\;+\frac{\delta}{2}\int_{\Omega}\psi|\alpha\nabla\varphi_{\delta}-D\nabla\psi_{\delta}|^{2}\,{\rm d}{\bm{x}}+\int_{\Omega}\frac{|{\bm{x}}|^{2}}{2}\psi_{\delta}R_{0}\,{\rm d}{\bm{x}}\,.$
We now can apply the Gronwall lemma, thanks to (6), (13) and the new bound (30), and conclude (10).
Step 2. The bounds in Proposition 2. The proof of the stronger gradient control (18) now needs to also rely on the Gronwall lemma. In fact the entropy has an extra term on the right-hand-side of (22) which can be bounded by
$\delta\int_{\Omega}\nabla\psi_{\delta}(\alpha\nabla\varphi_{\delta}-D\nabla\psi_{\delta})\,{\rm d}{\bm{x}}=$ $-\delta D\int_{\Omega}|\nabla\psi_{\delta}|^{2}\,{\rm d}{\bm{x}}+\int_{\Omega}\left(\nabla\psi_{\delta}\delta^{\frac{1}{2}}D^{\frac{1}{2}}\right)\left(\frac{\delta^{\frac{1}{2}}}{D^{\frac{1}{2}}}\alpha\nabla\varphi_{\delta}\right)\,{\rm d}{\bm{x}}$
$\leq$ $-\delta\frac{D}{2}\int_{\Omega}|\nabla\psi_{\delta}|^{2}\,{\rm d}{\bm{x}}+\delta\frac{\alpha}{2D}\int_{\Omega}|\nabla\varphi|^{2}\,{\rm d}{\bm{x}}\,.$
Then combining the entropy with $\partial_{t}\left[\frac{|\nabla\varphi|^{2}}{2}(1-\varphi)\right]$, the above bound and those detailed in Step 1 of Proposition 2 yield
$\frac{d}{dt}\int_{\Omega}\bigg[\frac{\alpha}{\gamma}\frac{|\nabla\varphi|^{2}}{2}(1-\varphi)+\psi\ln(\psi)\bigg]\,{\rm d}{\bm{x}}\leq\delta\frac{\alpha}{2D}\int_{\Omega}|\nabla\varphi|^{2}\,{\rm d}{\bm{x}}+R_{0}\int_{\Omega}\psi\,{\rm d}{\bm{x}}+R_{0}\int_{\Omega}\psi|\ln(\psi)|\,{\rm d}{\bm{x}}\,.$
From here we can apply the Gronwall lemma thanks to bounds (9), (11) and (7), concluding (23). Then bound (18) is obtained from the same calculation, as outlined in Step 1 of Proposition1, knowing (23). The proof of bounds (17), (19)-(21) remains unchanged.
Step 3. Local compactness. The proof of local compactness follows the same steps as in Section 4, with space compactness and bound (24) retrieved analogously. To gain time-compactness of $\varphi^{\frac{1}{2}}_{\delta}\psi_{\delta}$ we need to bound extra terms appearing in (25) which reads as
$\partial_{t}(\varphi_{\delta}^{\frac{1}{2}}\psi_{\delta})=$ $-\nabla\cdot\left[\,\varphi_{\delta}^{\frac{1}{2}}\psi_{\delta}\,{{\bm{V}}}(\psi_{\delta},\varphi_{\delta})\,\right]+\;\psi_{\delta}\nabla\varphi_{\delta}^{\frac{1}{2}}\cdot{{\bm{V}}}(\psi_{\delta},\varphi_{\delta})+\psi_{\delta}\varphi_{\delta}^{\frac{1}{2}}R(\psi_{\delta},\varphi_{\delta})-\frac{\gamma}{2}\varphi_{\delta}^{\frac{1}{2}}\psi_{\delta}^{2}$
$-\nabla\cdot\left[\,\delta\,\varphi_{\delta}^{\frac{1}{2}}\psi_{\delta}\,\left(\alpha\nabla\varphi_{\delta}-D\nabla\psi_{\delta}\right)\,\right]+\,\delta\,\psi_{\delta}\nabla\varphi_{\delta}^{\frac{1}{2}}\cdot\left(\alpha\nabla\varphi_{\delta}-D\nabla\psi_{\delta}\right)\,.$
Using the Cauchy-Schwarz inequality we see that
$\delta\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\varphi_{\delta}^{\frac{1}{2}}\psi_{\delta}\,|\alpha\nabla\varphi_{\delta}-D\nabla\psi_{\delta}|\,{\rm d}{\bm{x}}{\rm d}t\leq\delta\left[\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi_{\delta}\,{\rm d}{\bm{x}}{\rm d}t\;\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\varphi_{\delta}\psi_{\delta}|\alpha\nabla\varphi_{\delta}-D\nabla\psi_{\delta}|^{2}\,{\rm d}{\bm{x}}{\rm d}t\right]^{\frac{1}{2}}$
is bounded in $L^{1}_{\text{loc}}((0,T),\Omega)$, as the first integral is bounded by (9) and the second can be bounded by (13) knowing (8) under assumption (7). Similarly, we have that
$\delta\int_{0}^{T}\hskip-5.0pt\int_{\Omega}|\nabla\varphi_{\delta}^{\frac{1}{2}}\cdot\psi_{\delta}\left(\alpha\nabla\varphi_{\delta}-D\nabla\psi_{\delta}\right)|\,{\rm d}{\bm{x}}{\rm d}t\leq\delta^{\frac{1}{2}}\left[\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi_{\delta}|\nabla\sqrt{\psi_{\delta}}|^{2}\,{\rm d}{\bm{x}}{\rm d}t\;\delta\int_{0}^{T}\hskip-5.0pt\int_{\Omega}\psi_{\delta}|\alpha\nabla\varphi_{\delta}-D\nabla\psi_{\delta}|^{2}\,{\rm d}{\bm{x}}{\rm d}t\right]^{\frac{1}{2}}$
is bounded in $L^{1}_{\text{loc}}((0,T),\Omega)$, since the first integral is bounded by (18) and the second by the new bound (30). Then we can apply the Aubins-Lions-Simons lemma, concluding the proof.
6.2 Construction of solutions
The non-degenerate problem (29), still highly nonlinear, can be handled following the arguments in [35, 36, 37] and existence of global strong solutions follows. The solutions satisfy the properties of Proposition 5 uniformly in $\delta$.
Theorem 6 (Global existence of solutions)
Let $\Omega\equiv\mathbb{R}^{d}$ $(\psi_{\delta},\varphi_{\delta})$ denote a global solution of the non-degenerate problem (29) complemented with $\psi_{\delta}(0,{\bm{x}})=\psi^{0}_{\delta}$ and $\varphi_{\delta}(0,{\bm{x}})=\varphi^{0}_{\delta}$ satisfying the properties (6) and (7). Then, as $\delta\to 0$, $(\psi_{\delta},\varphi_{\delta})$ admits a subsequence converging to the pair $(\psi,\varphi)$, a weak solution of the degenerate problem (1)-(5) in the sense of Definition 3.
Proof. Taking for granted by [35, 36, 37] that there exists a global solution of the non-degenerate problem (29), under assumptions (6) and (7), we know that it satisfies all properties of Proposition 5. Then, since all properties are satisfied uniformly in $\delta$, letting $\delta\to 0$ we can pass to the limit along with subsequences similarly to as done in Section 5, and the limit pair $(\psi,\varphi)$ satisfies all properties of Propositions 1 and 2. This shows the global existence of weak solutions of the degenerate problem (1).